The lateral edge of a square prism makes an angle of 60º with the base. What is the altitude of the prism if the lateral edge is √3cm


1
Expert's answer
2021-02-16T11:13:47-0500

"h(altitude) = sin60\degree\cdot\sqrt{3} = \sqrt{3}/2 \cdot \sqrt{3} = 3/2"